# Wasserstein Convergence of Score-based Generative Models under Semiconvexity and Discontinuous Gradients

### v1

@@ 3 Wasserstein Convergence Analysis for SGMs
In this section, we provide the full non-asymptotic estimates in Wasserstein distance of order two between the target data distribution $\pi_{\mathsf{D}}$ and the generative distribution of the diffusion model under the assumptions stated below. As explained in [7, Section 2 and Appendix A]), it could be necessary to restrict $t\in[\epsilon,T]$ for $\epsilon\in(0,1)$ in (5) to account for numerical instability issues for training and sampling at $t=0$ as observed in practice in [57, Appendix C] and for the possibility that the integral of the score function in (5) may diverge when $t=0$. Therefore, we truncate the integration in the backward diffusion at $T-\epsilon$ and run the process $(Y_{t})_{t\in[0,T-\epsilon]}$.
@@ 3.1 Assumptions
We begin by stating the main assumptions of our setting. The optimization problem in (5) can be solved using algorithms such as ADAM [33], stochastic gradient descent [29], Stochastic Gradient Langevin Dynamics [7, Section 3.1], and TheoPouLa [43], provided they satisfy the following assumption.
@@
Let $\theta^{*}$ be a minimiser [^1] The score-matching optimization problem (5) is not necessarily (strongly) convex. of (5) and let $\hat{\theta}$ be the (random) estimator of $\theta^{*}$ obtained through some approximation procedure such that $\mathbb{E}[|\hat{\theta}|^{2}]<\infty$. There exists $\widetilde{\varepsilon}_{\text{AL}}>0$ such that
@@
Recall that for any $x\in\mathbb{R}^{d}$, $h(x)$ denotes an element of the subdifferential $\partial U(x)$ defined in (1).
@@
    The potential $U$ is $K$-semiconvex. That is, there exists $K,R\geq 0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
@@
    The potential $U$ is $\mu$-strongly convex at infinity. That is, there exists $\mu>0$ and $R\geq 0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq\mu|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\geq R.$ (9)
Remark 3.2.
As a consequence of Proposition A.1, due to in [15, Proposition 3.1], and Assumption 2-(i), $\nabla\log p_{t}(x)$ is continuous for $t\in(0,T]$ and $x\in\mathbb{R}^{d}$. Moreover, Assumption 2 implies that the processes in (3) and (4) have a unique strong solution.
Next, we consider the following assumption on the approximating function $s$.
@@
where $D_{1}$, $D_{2}$ and $D_{3}$ have the following growth in each variable: i.e. there exist $\mathsf{K}_{1}$, $\mathsf{K}_{2}$, and $\mathsf{K}_{3}>0$ such that for each $t,\bar{t}\in[0,T]$ and $\theta,\bar{\theta}\in\mathbb{R}^{M}$,
@@
Remark 3.3.
Assumption 3.a implies that the process in (6), (7), and (8) have a unique strong solution. We refer to [7, Remark 6] for discussion regarding the practical justification of this assumption when it is applied to neural networks.
By adding a further condition on the gradient of $s$ in Assumption 3.a we are able to achieve the optimal rate of convergence in Theorem 3.14 below.
@@
For the following assumption on the score approximation, we let $\hat{\theta}$ be as in Assumption (1) and we let $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ be the auxiliary process defined in (6).
@@
Remark 3.4.
We refer to [7, Remark 7, 8, and 9] for discussion regarding the theoretical and practical soundedness of Assumption 4, which is a now a standard assumption considered in the literature, see e.g. [7, 20, 21, 58].
@@ 3.2 Assumption 2 and Weak Convexity of the Data Distribution
We show that Assumption 2-(ii) and Assumption 2-(iii) are related to the notion of weak convexity in the sense made precise in Proposition 3.7 below. We start by introducing the definition of weak convexity for subgradients.
Definition 3.5.
@@
$\kappa_{U}(r)=\inf_{x,\bar{x}\in\mathbb{R}^{M}:\ |x-\bar{x}|=r}\left\{\frac{\langle\partial U(x)-\partial U(\bar{x}),x-\bar{x}\rangle}{|x-\bar{x}|^{2}}\right\}$ (11)
satisfies
$\kappa_{U}(r)\geq\beta-r^{-1}f_{L}(r),\quad\text{for all}\ \ r>0.$ (12)
@@
[…]
Lemma 3.6.
[16, Modification of Lemma 5.9] Assume that U is weakly convex as in Definition 3.5 and fix $t\in[0,T]$. Then, the function $x\mapsto-\log p_{t}(x)$ is weakly convex with weak convexity profile $\kappa_{-\log p_{t}(x)}$ satisfying
@@
$\langle\nabla\log p_{t}(x)-\nabla\log p_{t}(\bar{x}),x-\bar{x}\rangle\leq-\widehat{C}_{t}|x-\bar{x}|^{2},\quad\text{for}\ x,\bar{x}\in\mathbb{R}^{\text{d}}$ (14)
@@
An overview of the proof of Proposition 3.7 below can be found in Appendix B.
Proposition 3.7.
Let the data distribution $\pi_{\mathsf{D}}$ be in Assumption 2 and let $f_{L}$ be as in (13). Then
@@
where $\mu>0$ is the strong convexity at infinity constant from Assumption 2-(iii). Conversely, if $U$ is weakly convex as in Definition 3.5 with lower bound (16) for some known constants $\mu$ and $L>0$, then
@@
    The potential $U$ is $\widetilde{\mu}$-strongly convex at infinity with $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)>0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$, we have
@@
    which holds for all $R>0$ when $\mu>L$ and for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$ with $z_{0}$ being the solution of (43) when $\mu\leq L$.
@@
    The potential $U$ is $K$-semiconvex, such that there exists $K,R\geq 0$ for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq-K|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\leq R.$ (18)
As a consequence of Proposition 3.7 and Lemma 3.6, one obtains the explicit form of $\widehat{C}_{t}$ in (14) in our setting, which is given in the following corollary.
Corollary 3.8.
Let $U$ be $K$-semiconvex as in Assumption 2-(ii) and be $\mu$-strongly convex at infinity as in Assumption 2-(iii) and fix $t\in[0,T]$. Then
@@
for some $L>0$ satisfying (16).
Remark 3.9.
By Corollary 3.8 and the proof of Proposition 3.7, we have
$\lim_{t\rightarrow 0}\beta^{\text{OS}}_{t}=\mu-L<-K.$ (21)
@@
[…]
For this reason, we use the constant $K+\mu$ as a proxy of the constant $L$ and replace (20) with the following monotonicity bound
@@
which is consistent with $\pi_{\infty}\sim\mathscr{N}(0,I_{d})$, the invariant distribution of the OU process.
Using the explicit expression of (22), we are able to find a time for which the integral of the monotonicity bound [^3] Note that $\beta^{\text{OS},K,\mu}_{t}$ is a function of time. $\beta^{\text{OS},K,\mu}_{t}$ is positive. The proof of the following result is postponed to Appendix B.
Proposition 3.10.
Let $\mu>0$ and $K\geq 0$. The time integral of $\beta^{\text{OS},K,\mu}_{t}$ from Remark 3.9 is
@@
when $t>t^{\star}>\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ with $t^{\star}:=\inf\left\{t>0:B(t,0,\mu,K)>0\right\}$.
Remark 3.11.
If we consider the case when $K=0$ in Assumption 2-(ii), then (23) is satisfied for all $t>0$.

### v2

@@ 3 Wasserstein Convergence Analysis for SGMs
In this section, we provide the full non-asymptotic estimates in Wasserstein distance of order two between the target data distribution $\pi_{\mathsf{D}}$ and the generative distribution of the diffusion model under the assumptions stated below. As discussed in Bruno et al. (2025, Section 2 and Appendix A), it may be necessary to restrict $t\in[\epsilon,T]$ for $\epsilon\in(0,1)$ in 4 to account for numerical instabilities that can arise during training and sampling near $t=0$ as also observed in practice in Song et al. (2021, Appendix C), and for the possibility that the integral of the score function in 4 may diverge when $t=0$. Therefore, we truncate the integration in the backward diffusion at $T-\epsilon$ and consider the process $(Y_{t})_{t\in[0,T-\epsilon]}$.
@@ 3.1 Assumptions
We begin by stating the main assumptions of our setting. The optimization problem in 4 can be solved using algorithms such as stochastic gradient descent (Jentzen et al., 2021), ADAM (Kingma & Ba, 2015), Stochastic Gradient Langevin Dynamics (Bruno et al., 2025, Section 3.1), and TheoPouLa (Lim & Sabanis, 2024), provided they satisfy the following assumption.
@@
Let $\theta^{*}$ be a minimiser [^3] The score-matching optimization problem 4 is not necessarily (strongly) convex. of 4 and let $\hat{\theta}$ be the (random) estimator of $\theta^{*}$ obtained through some approximation procedure such that $\mathbb{E}[|\hat{\theta}|^{2}]<\infty$. There exists $\widetilde{\varepsilon}_{\text{AL}}>0$ such that
@@
In this work, we consider the potentials to be semiconvex functions– a broad generalization of convex functions that includes non-convex functions whose curvature is bounded from below. This class allows for discontinuities in the gradient while retaining key analytical properties of convex functions, such as the existence of well-defined subgradients.
Definition 2.
A function $U$ is semiconvex if there exists $K\geq 0$ such that $U+\frac{K}{2}|\cdot|^{2}$ is convex.
Semiconvexity has received significant attention [^4] Some of these references refer to semiconvex functions as weakly convex. We avoid this terminology to prevent confusion with the notion of weak convexity introduced in Definition 10 below. in the machine learning community (Davis et al., 2018; Sun & Yu, 2019; Richards & Rabbat, 2021; Liu et al., 2021; Rafique et al., 2022), optimization (Li et al., 2020; Ma et al., 2020; Li & Xu, 2021; Hu et al., 2025), optimal control (Cannarsa & Sinestrari, 2004), and the study of fully nonlinear partial differential equations (Braga et al., 2019; Payne & Redaelli, 2023). We refer the reader to Duda & Zajıcek (2009); Cattiaux & Guillin (2014) for comprehensive overviews of the mathematical challenges associated with this class. Importantly, semiconvex functions may admit discontinuous gradients which are characterized using the Fréchet subdifferential (Bazaraa et al., 1974; Alberti et al., 1992).
Definition 3.
For a function $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define the subdifferential $\partial U(x)$ of $U$ at $x\in\mathbb{R}^{d}$ as
$\partial U(x)=\left\{\tilde{p}\in\mathbb{R}^{d}:\ \liminf_{z\rightarrow x}\frac{U(z)-U(x)-\langle\tilde{p},z-x\rangle}{|z-x|}\geq 0\right\}.$ (8)
The set 8 is closed and convex, and may be empty in general. We say that $U$ is Fréchet subdifferentiable at $x$ if $\partial U(x)\neq\emptyset$. Any element $h(x)\in\partial U(x)$ is called a Fréchet subgradient of $U$ at $x\in\mathbb{R}^{d}$. When $U$ is differentiable at $x$, the subdifferential reduces to $\partial U(x)=\{\nabla U(x)\}$. Crucially, semiconvex functions are Fréchet subdifferentiable at every points in $\mathbb{R}^{d}$, i.e. $\partial U(x)\neq\emptyset$ for all $x$ (Alberti et al., 1992; Cannarsa & Sinestrari, 2004). In this case, $h(x)+Kx$ corresponds to the classical convex subgradient (Vial, 1983, Proposition 4.6). Moreover, every element of the subdifferential of a semiconvex function satisfies a one-sided Lipschitz condition, ensuring the existence of $h(x)\in\partial U(x)$. The following lemma– adapted from (Alberti et al., 1992, Proposition 2.1) and presented in (Johnston et al., 2025, Lemma 1)–and its subsequent corollary formalize this property.
Lemma 4.
(Alberti et al., 1992, Modification of Proposition 2.1) Let $U$ be a semiconvex function. Then, $U$ is locally Lipschitz continuous, the subdifferential set $\partial U$ is non-empty, compact, and $\tilde{p}\in\partial U(x)$, if and only if
$U(z)-U(x)-\langle\tilde{p},z-x\rangle\geq-\frac{K}{2}|z-x|^{2},$
for all $x,z\in\mathbb{R}^{d}$.
Corollary 5.
(Johnston et al., 2025, Corollary 1) Let $x,z\in\mathbb{R}^{d}$, $\tilde{p}\in\partial U(x)$, and $\tilde{q}\in\partial U(z)$. Then,
$\langle\tilde{p}-\tilde{q},x-z\rangle\geq-K|x-z|^{2}.$
We state the assumption on the target data distribution $\pi_{\mathsf{D}}$ below. Recall that $h(x)\in\partial U(x)$ is the Fréchet subgradient of $U$ at $x\in\mathbb{R}^{d}$.
@@
    The potential $U$ is $K$-semiconvex (on a ball). That is, there exists $K,R\geq 0$, such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
@@
    The potential $U$ is $\mu$-strongly convex at infinity [^5] Intuitively, outside a sufficiently large region, $U$ bends upwards at least as much as a quadratic function.. That is, there exists $\mu>0$ such that for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq\mu|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|\geq R,$ (9)
    where $R$ is the same as in point (ii).
Remark 6.
As a consequence of Proposition 23, due to Conforti et al. (2025, Proposition 3.1), and Assumption 2-(i), we have that, for any $t\in(0,T)$, the map $x\mapsto\nabla p_{t}(x)$ is continuously differentiable, and for any $x\in\mathbb{R}^{d}$, the map $t\mapsto p_{t}(x)$ is continuously differentiable on $(0,T]$. Moreover, Assumption 2 implies that the processes in 2 and 3 have a unique strong solution.
Next, we consider the following assumption on the approximating function $s$, which is also adopted in Bruno et al. (2025, Assumption 3.a).
@@
where $D_{1}$, $D_{2}$ and $D_{3}$ have the following growth in each variable: i.e., there exist $\mathsf{K}_{1}$, $\mathsf{K}_{2}$, and $\mathsf{K}_{3}>0$ such that for each $t,\bar{t}\in[0,T]$ and $\theta,\bar{\theta}\in\mathbb{R}^{M}$,
@@
Remark 7.
Assumption 3.a requires that the approximating function $s$ is Lipschitz continuous in both the input variable $x$ and the parameter $\theta$. In time $t$, it allows $s$ to be either Hölder continuous (for $\alpha\in[\frac{1}{2},1)$) or Lipschitz continuous (for $\alpha=1$). This relaxed continuity in $t$ for the drift term of 7 is standard for the Euler–Maruyama schemes for SDEs. Crucially, we show in Theorems 19 and 21 that this weaker condition in $t$ still guarantees convergence of the generative algorithm to $\pi_{\mathsf{D}}$. As noted by Bruno et al. (2025, Remark 6) in the context of neural network-based approximations, Assumption 3.a, when $\alpha=1$, covers the case where $s$ is implemented as a neural network with a hyperbolic tangent or sigmoid activation function at the final layer. Moreover, Assumption 3.a implies that the process in 5, 6, and 7 have a unique strong solution.
Remark 8.
Let $\mathsf{K}_{\text{Total}}:=\mathsf{K}_{1}+\mathsf{K}_{2}+\mathsf{K}_{3}+|s(0,0,0)|>0$. Using Assumption 3.a, one obtains
$\begin{split}|s(t,\theta,x)|&\leq\mathsf{K}_{\text{Total}}(1+|t|^{\alpha})(1+|\theta|+|x|).\end{split}$
The proof of Remark 8 can be found, e.g., in Bruno et al. (2025, Appendix D.3). By imposing an additional condition on the gradient of $s$ in Assumption 3.a—as done in Bruno et al. (2025, Assumption 3.b)—, we obtain the optimal convergence rate established in Theorem 21 below.
@@
For the following assumption on the score approximation, we let $\hat{\theta}$ be as in Assumption 1 and we let $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ be the auxiliary process defined in 5.
@@
Remark 9.
Assumption 4 is now a standard assumption considered in the literature, see, e.g., Gao et al. (2025); Bruno et al. (2025); Strasman et al. (2025); Silveri & Ocello (2025), and its theoretical and practical soundedness is discussed, e.g., in Bruno et al. (2025, Remark 7, 8, and 9).
@@ 3.2 Assumption 2 and Weak Convexity of the Data Distribution
We start by introducing the definition of weak convexity, a concept that has been widely used in conjunction with coupling techniques to analyze the long-time behavior of gradient flows SDEs (Conforti, 2023; 2024; Conforti et al., 2023b; a) [^6] Recently, this notion has been used in the context of score-based generative models in Silveri & Ocello (2025, Definition 3.1)., and we extend its application here to subgradients of $U$.
Definition 10.
@@
$\kappa_{U}(r)=\inf_{x,\bar{x}\in\mathbb{R}^{d}:\ |x-\bar{x}|=r}\left\{\frac{\langle h(x)-h(\bar{x}),x-\bar{x}\rangle}{|x-\bar{x}|^{2}}\right\},$ (11)
where $h(x)\in\partial U(x)$ is the subgradient of $U$ at $x\in\mathbb{R}^{d}$, satisfies
$\kappa_{U}(r)\geq\beta-r^{-1}f_{L}(r),\quad\text{for all}\ \ r>0,$ (12)
@@
Remark 11.
The weak convexity profile $\kappa_{U}$ in 11 serves as an averaged/integrated convexity lower bound for the potential $U$, evaluated over pairs of points separated by a distance $r>0$. Unlike the standard convexity condition $\kappa_{U}\geq 0$, which characterizes convex (or log-concave) potentials, Definition 10 allows $\kappa_{U}$ to vary with $r$, thereby admitting non-uniform lower bounds. This generalization yields a broader and more flexible notion of convexity that extends beyond classical log-concavity, making Definition 10 substantially weaker than log-concavity (Conforti, 2023; 2024; Conforti et al., 2023b; a; Silveri & Ocello, 2025).
We modify Conforti et al. (2023b, Lemma 5.9) to our setting, namely when $\beta>0$ [^7] See Silveri & Ocello (2025, Lemma B.4) for a similar statement. to have an explicit expression of the weak convexity constant at each $t\in(0,T]$.
Lemma 12.
(Conforti et al., 2023b, Modification of Lemma 5.9) Assume that U is weakly convex as in Definition 10 and fix $t\in(0,T]$. Then, the function $x\mapsto-\log p_{t}(x)$ is weakly convex with weak convexity profile $\kappa_{-\log p_{t}(x)}$ satisfying
@@
$\langle\nabla\log p_{t}(x)-\nabla\log p_{t}(\bar{x}),x-\bar{x}\rangle\leq-\widehat{C}_{t}|x-\bar{x}|^{2},\quad\text{for}\ x,\bar{x}\in\mathbb{R}^{\text{d}},$ (14)
@@
We show that Assumption 2-(ii) and Assumption 2-(iii) are related to the notion of weak convexity (Definition 10) in the sense made precise in Proposition 13 below. An overview of the proof of Proposition 13 below can be found in Appendix B.
Proposition 13.
Let the data distribution $\pi_{\mathsf{D}}$ be in Assumption 2, and let $f_{L}$ and $L>0$ be as in Definition 10. Then the potential $U$ is weakly convex as in Definition 10 with
@@
where $\mu>0$ in 16 is the strong convexity at infinity constant from Assumption 2-(iii). Conversely, if $U$ is weakly convex as in Definition 10 with lower bound 16 for some known constants $\mu$ and $L>0$, then
@@
    The potential $U$ is $\widetilde{\mu}-$strongly convex at infinity with $\widetilde{\mu}:=\mu-R^{-1}f_{L}(R)>0$ [^8] We refer to the proof of Proposition 13 in Appendix B below for the derivation of this constant., such that for all $x,\bar{x}\in\mathbb{R}^{d}$, we have
@@
    which holds for all $R>0$ when $\mu>L$ and for $R\geq R_{0}=\frac{2z_{0}}{L^{1/2}}$ with $z_{0}$ being the solution of 45 when $\mu\leq L$.
@@
    The potential $U$ is $K$-semiconvex, such that there exists $K\geq 0$ for all $x,\bar{x}\in\mathbb{R}^{d}$,
    $\langle h(x)-h(\bar{x}),x-\bar{x}\rangle\geq-K|x-\bar{x}|^{2},\qquad\text{when}\quad|x-\bar{x}|<R,$ (18)
    where $R$ is the same as in point (2).
As a consequence of Proposition 13 and Lemma 12, one obtains the explicit form of $\widehat{C}_{t}$ in 14 in our setting, which is given in the following corollary.
Corollary 14.
Let $U$ be $K$-semiconvex as in Assumption 2-(ii) and be $\mu$-strongly convex at infinity as in Assumption 2-(iii) and fix $t\in(0,T]$. Then
@@
for some $L>0$ satisfying 16.
Remark 15.
By Corollary 14 and the proof of Proposition 13, we have
$\lim_{t\rightarrow 0}\beta^{\text{OS}}_{t}=\mu-L<-K,$ (21)
which shows that $\beta^{\text{OS}}_{t}$ is not the lowest bound for the left-hand side of 19. We emphasize that the gap between the limit on the left-hand side of 21 and the semiconvexity constant $K$ is due to the particular choice of $f_{L}$ in 13 in Proposition 13. This gap may vanish if we replace $f_{L}$ with an appropriate function $f\in\mathcal$$\widetilde${F}}(Conforti et al., 2023a, Section 2.1.2), (Conforti et al., 2023b, Section 5.3.1), where
@@
Note that $\mathcal$$\widetilde${F}} is non-empty and contains $r\mapsto 2\tanh(r/2)$. For this reason, we use the constant $K+\mu$ as a proxy of the constant $L$ and replace 20 with the following monotonicity bound
@@
which is consistent with $\pi_{\infty}\sim\mathcal{N}(0,I_{d})$, the invariant distribution of the OU process.
Using the explicit expression of 22, we are able to find a time for which the integral of the monotonicity bound [^10] Note that $\beta^{\text{OS},K,\mu}_{t}$ is a function of time. $\beta^{\text{OS},K,\mu}_{t}$ is positive. The proof of the following result is postponed to Appendix B.
Proposition 16.
Let $\mu>0$ and $K\geq 0$. The time integral of $\beta^{\text{OS},K,\mu}_{t}$ from Remark 15 is
@@
when $t>t^{\star}>\bar{t}:=\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)$ with $t^{\star}:=\inf\left\{t>0:B(t,0,\mu,K)>0\right\}$.
Remark 17.
If we consider the case when $K=0$ in Assumption 2-(ii), then 23 is satisfied for all $t>0$.
Figure 1: Score function 25 for fixed $x$ and time $\bar{t}$.
Remark 18.
We provide a numerical illustration of the critical time of Proposition 23 when a non-log-concave distribution becomes log-concave in the case when $\pi_{\mathsf{D}}$ is a one-dimensional Gaussian mixture with two equi-weighted modes $\eta=2$, each mode having same variance $s^{2}=9$, namely
$\begin{split}\pi_{\mathsf{D}}(\mathrm{d}x)&=(2(18\pi)^{1/2})^{-1}\left(\exp\left(-\frac{|x-2|^{2}}{18}\right)+\exp\left(-\frac{|x+2|^{2}}{18}\right)\right)\mathrm{d}x.\end{split}$ (24)
For this choice of $\pi_{\mathsf{D}}$, the semiconvexity constant $K=\frac{2\eta^{2}}{(s^{2})^{2}}=\frac{8}{81}$, and the strong convexity at infinity constant $\mu=\frac{s^{2}-2\eta^{2}}{(s^{2})^{2}}=\frac{1}{81}$. Here, $R\geq 0$. The score function of 24 is given by
$\begin{split}\nabla\log p_{t}(x)=-\frac{x}{9m_{t}^{2}+\sigma_{t}^{2}}+\frac{2m_{t}}{9m_{t}^{2}+\sigma_{t}^{2}}\frac{\exp\left\{\frac{-\left(x-2m_{t}\right)^{2}}{2(9m_{t}^{2}+\sigma_{t}^{2})}\right\}-\exp\left\{\frac{-\left(x+2m_{t}\right)^{2}}{2(9m_{t}^{2}+\sigma_{t}^{2})}\right\}}{\exp\left\{\frac{-\left(x-2m_{t}\right)^{2}}{2(9m_{t}^{2}+\sigma_{t}^{2})}\right\}+\exp\left\{\frac{-\left(x+2m_{t}\right)^{2}}{2(9m_{t}^{2}+\sigma_{t}^{2})}\right\}},\quad t\in[0,T],\ x\in\mathbb{R},\end{split}$ (25)
where $m_{t}=e^{-t}$ and $\sigma_{t}^{2}=1-e^{-2t}$ comes from the representation of the OU process in 52. Figure 1 displays the time behaviour of score function 25 over the time interval $[0,10]$ for fixed values x=0.8x=-0.8 and $x=0.5$. We also indicate the time $\bar{t}=\ln\left(\sqrt{1+\frac{K}{\mu^{2}}}\right)\approx 3.2377<t^{\star}$ from Proposition 25. The score function in Figure 1 converges to $\nabla\log(\pi_{\infty}(x))=-x$, where $\pi_{\infty}=\mathcal{N}(0,1)$.